Analyze whether to invest or not in a riskless asset whose price is 33000, the certain stream of revenue is 2200euros per year and the annual interest rate on government bond is 3%. Make explicit your assumptions
1
Expert's answer
2014-05-27T08:30:36-0400
P = 33000, TR = 2200 euros per year, annual interest rate on government bond is 3%.
If we earn 2200 euros per year after investment of 33000 euros, our annual interest rate of this investment is r = 2200/33000 = 0.067 or 6.7%

As annual interest rate on government bond is 3%, which also has no risks, the proposed investment is better for us than investing in government bonds, as from the bonds we can earn only TR = 33000*0.03 = 990 euros per year.'